OpenStudy (anonymous):

Perimeter of the equilateral triangle ABC is 72 and the perimeter of the isosceles triangle DBC with the base BC=24 is 50. Which is AB + BD?

OpenStudy (anonymous):

AB + BD = 13 + 24 = 37 Each side of the equilateral triangle ABC is 72/3 = 24 in length. For the isosceles triangle: 2 x + 24 = 50 where x is one of the equal sides. x = 13